OpenStudy (anonymous):

Solve for x. 3x+1/14=2x/8

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

how and why?

OpenStudy (anonymous):

give me a minute i can explain perfactly

OpenStudy (anonymous):

Hey! Is 3x + 1 in the numerator or is it just 1 in the numerator?

OpenStudy (anonymous):

one sec.

OpenStudy (anonymous):

der

OpenStudy (anonymous):

Ah, ok! Do you know how to cross multiply?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That's what we need to do!

OpenStudy (anonymous):

ow this is a strange way of doing it that way

OpenStudy (anonymous):

how*

OpenStudy (anonymous):



OpenStudy (anonymous):

\(\bf\Large \frac{a}{b} = \frac{c}{d} \\\\~ \\\\\to ad=bc\)

OpenStudy (anonymous):

is this right

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

icri evry teim :(

OpenStudy (anonymous):

After cross multiplying, we get: 8(3x+1) = 14(2x) Do you know the distributive property?

OpenStudy (anonymous):

12x+4=7x2x?

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

okay then :/

OpenStudy (anonymous):

The distributive property is a(b+c) = ab + ac So, we have 8(3x + 1) What is 8 * 3x = ? Remember, x is just a variable and what is 8 * 1 = ?